# Reverberation mapping of high-mass and high-redshift quasars using gravitational time delays

### v1

@@ Abstract
Mass estimates of black holes (BHs) in the centers of Active Galactic Nuclei (AGN) often rely on the radius-luminosity relation. However, this relation, usually probed by reverberation mapping (RM), is poorly constrained in the high-luminosity and high-redshift ends due to the very long expected lag times. Multiply imaged AGN may thus offer a unique opportunity to explore the radius-luminosity relation at these ends. In addition to comprising several magnified images which enable a more efficient light-curve sampling, the time delay between multiple images of strongly lensed quasars can also aid in making such RM measurements feasible on reasonable timescales: If the time delay is, for example, of the order of the expected time lag, changes in the emission lines in the leading image can be observed around the same time as the changes in the continuum in the trailing image. In this work we probe the typical time-delay distribution in galaxy-cluster lenses and estimate the number of both high-mass ( M⊙), and high-redshift () quasars that are expected to be strongly lensed by clusters. We find that less than one, very massive (and luminous, L erg s-1) multiply imaged quasar is expected across the sky down to 30 AB mag. Nonetheless, up to several dozen thousand M- M⊙ broad-line AGN at should be multiply imaged by galaxy clusters and detectable with JWST, hundreds with Euclid and several thousands with the Roman Space Telescope, across the whole sky. These could supply an important calibration for the BH mass scaling in the early Universe.
@@ 1 Introduction
Super massive black holes (SMBHs) are found in the centers of most galaxies. The masses of those SMBHs are of much interest as their evolution is not well understood, and since they are believed to play a major role in their host galaxy evolution. In fact, tight relations seem to hold between the mass of an SMBH and the properties of its host galaxy such as the luminosity or the velocity dispersion of the bulge (e.g., Ferrarese & Merritt, 2000; Kormendy & Ho, 2013; Reines & Volonteri, 2015; Ding et al., 2020). Measuring SMBH masses as a function of redshift can also shed light onto black hole (BH) formation and the physics of accretion processes (e.g., Wang et al., 2021; Volonteri, 2012), and their mass evolution through cosmological time might help constrain different cosmological models (e.g., Fanidakis et al., 2012; Habouzit et al., 2022).
@@
Unfortunately, RM of high-luminosity and high-redshift quasars faces several difficulties. High-luminosity quasars are presumed to have large BLR sizes, resulting in long time lags between continuum and emission-line variations. Similarly, high-redshift quasars and AGN (e.g., Wang et al., 2021; Furtak et al., 2024; Larson et al., 2023; Goulding et al., 2023) will also show longer observed time lags compared to their lower-redshift counterparts, due to cosmological time dilation. As a result, observations for several decades are needed for those RM campaigns (Kaspi, 2018, 2021). Nevertheless, several attempts have been recently made to conduct RM for those types of AGN. In one study, Kaspi et al. (2021) observed photometrically 11 quasars at redshifts and luminosities for almost twenty years while six of them were monitored spectrophotometrically for 13 years. Their study manages to significantly measure time lags of the C iv Å doublet in three objects and a C iii] Å time lag in one object. Scaling the BLR size with the UV continuum luminosity they conclude that their objects lie on the previously measured relation. Shen et al. (2023) preformed RM using 11-year photometric and a 7-year spectroscopic data for 849 broad-line quasars. They report 339 significantly measured time lag detections for the different emission lines. Although they claim that on average the R-L relations follow existing ones, they report an intrinsic scatter of dex in the lags for the H line at 4861 Å and the Mg ii] Å doublet, whereas the C iv R-L relation has an even more substantial scatter of dex. It is thus becoming important to understand whether these large scatters stem from intrinsic physical processes at high luminosities or due to the fact that there is often a difficulty to significantly measure time lags for such objects given the monitoring times needed.
@@
In this work, we discuss the possibility to probe the R-L relation for the high-luminosity or high-redshift regimes using multiply imaged quasars. If the strong-lensing (SL) time delays between different multiple images of a lensed quasar are of the same order as the time lags, the expected changes in the emission lines can be observed in a leading multiple image around the same time as the changes in the continuum from a trailing multiple image, so that RM of these types of objects can be done with a reasonable and shorter monitoring time-span than for similar, un-lensed quasars (nominally shorter by one lag-time, but as we also discuss here in practice the gain will be larger, as a lensed quasar will be lensed into several magnified images, supplying a more efficient sampling of the light curve).
@@
About a handful of quasars and AGN, mostly around cosmic noon, are known to be multiply imaged by galaxy clusters (e.g., Inada et al., 2003; Oguri et al., 2013; Sharon et al., 2017; Acebron et al., 2022; Furtak et al., 2023a; Napier et al., 2023a), whereas some have also been used for cosmography (e.g., Napier et al., 2023b). Notably, a successful RM of a strongly lensed quasar has been recently done by Williams et al. (2021) where they exploit the time delays between the multiple images of the strongly lensed quasar SDSS J2222+2745 to construct a detailed light curve spending over 6 years in the observed frame with 4.5 years of observations with high cadence. They significantly detect a time lag of rest-frame days for the C iv emission line, whereas the time delays from the two other images they use are of and days. The quasar SDSS J2222+2745 ranged in continuum luminosity with the average value of log10[/erg s-1] = . In this illustrative case the time lag was short enough so that the gain from image multiplicity was mainly manifested in additional data points on the light curve, thus importantly allowing to shorten the overall monitoring time. Here, we wish to focus on objects at the high luminosity end, namely continuum UV luminosities erg s-1 (BH masses of log10M), or at high redshifts (i.e., ), for which we expect time lags of days, depending on redshift. For example, a redshift , M⊙ quasar (e.g. Wang et al., 2021) would be expected to show an observed time lag of about 3 or 6 years respectively, requiring decades of monitoring to measure it without lensing multiplicity. In this paper we suggest that SL time delays of lensed quasars may be a useful route to constraining the R-L relation at those regimes. This is becoming especially timely now that JWST has been uncovering a new abundant population of Type I AGN at high redshifts (e.g., Kocevski et al., 2023a; Harikane et al., 2023; Matthee et al., 2023; Larson et al., 2023; Greene et al., 2023), including the first example of a cluster-lensed multiply imaged one at 7 (Furtak et al., 2023b, 2024).
@@
The work is constructed as follows. In section 2 we start by estimating the expected time lags for high-luminosity and high-redshift quasars (§2.1), and then estimate the distribution of time delays between the multiple images of lensed systems in typical lensing clusters (§2.2). In section 3 we estimate the expected number of lensed quasars and AGN that could be found with e.g., the JWST, Euclid, or with the Roman Space Telescope, in the near future. In section 5 we demonstrate the suggested idea using simulated light curves, and in section 6 we conclude the work. Throughout we assume a standard flat CDM cosmology with , , and . Magnitudes are quoted in the AB system (Oke & Gunn, 1983) and all quoted uncertainties represent ranges unless otherwise stated.
@@
In this work we use the terms “quasars” and “AGN” synonymously. In addition, whenever we discuss lensed AGN in this work, we explicitly refer to multiply imaged examples.
2 The distributions of time lags and lensing time delays
2.1 RM of quasars: the distributions of time lags
@@
As was noted above, RM of quasars relies on the ability to detect temporal variations in the continuum emission from a quasar that lead to, after a time lag which depends on the BLR size, similar variations in the emission lines from the BLR. In this section, we estimate the BLR sizes – and thus time lags – of a representative sample of quasars, namely the SDSS quasar catalog, which is based on Data Release 7 (DR7) (Shen et al., 2011) and is available online. This catalog supplies continuum luminosities for either or , depending on redshift.
@@
The distributions of BLR sizes are shown in Fig.1. The expected H time-lag distribution is obtained assuming the R-L relation from Bentz et al. (2013), for those quasars that have a measurement of the continuum luminosity at 5100 Å. The expected C iv time-lag distribution is obtained assuming the R-L relation from Kaspi et al. (2021) (see their eq. 1), using the continuum UV luminosity at 1350 Å. Note that the C iv subsample consists of 52,427 objects, which have a higher redshift (median redshift of 1.96) and are typically brighter (median Å erg s-1) than the H subsample (23,093 objects, median redshift of 0.59, median Å erg s-1). As can be seen in Fig. 1, the lag-time distributions, plotted as log10(R) where R is in light days, are approximately Gaussian, with a mean and width of = and = for H and C iv, respectively. About %, roughly, of the quasars show C iv time lags of days or longer in the rest-frame, and about % of the quasars show H time lags of days or longer in the rest-frame. This translates to days for redshifts of , respectively, which we define here as the high redshift regime.
@@ 2.2 The distribution of strong-lensing time delays
In this subsection we aim to obtain a general distribution of the time delays caused by strong lensing in clusters. Galaxy clusters act as gravitational lenses, causing the bending of light. When the projected mass density is larger than some critical density for lensing, strong lensing will occur comprising multiple images of background sources. Using the thin lens approximation, light that travels from a source to the observer will be delayed by some time (Schneider, 1985):
@@
In order to create the time delay (TD) distribution, we randomly plant sources in the source plane and re-lens them to the image plane using the deflection fields from the model. Typically we use a 100 multiply imaged sources in each case. We calculate the time delays between all available pairs using eq. 1 and repeat this process for different source redshifts, in the ranges of , , , , and . The SL time-delay distributions are plotted in Fig.2.
@@
In Fig. 2 we also compare the lensing TD probability distributions to the time lags expected from the BLR sizes, recalculated to match each redshift bin. We see that there is an overlap in the distributions such that time delays by both types of galaxy clusters cover the entire range of expected time lags, with a higher probability of having the relevant time delays, i.e., of order the time lag, for higher-luminosity and higher-redshift quasars. About of the TD pairs in all clusters overlap with the range of time lags, depending on the redshift of the lensed source. We see that the peak of the TD distribution caused by lower mass clusters, which should be more abundant, agrees better with the peak of the time lag distributions, but those clusters would have also less multiple images per cluster in proportion to their smaller critical area size (which is roughly proportional to the square root of the mass), such that both populations should be useful. Also overplotted is a line marking a typical time delay expected by a massive galaxy lens – where we nominally adopt for this calculation a SIS galaxy of km s-1 placed at redshift 0.5 (Oguri, 2019), demonstrating that time delays in the case of lens galaxies are too short for a significant gain in the wait time needed for RM of high-redshift, high-mass quasars, and that galaxy clusters are needed for that purpose.
@@ 3 Expected numbers of lensed quasars
To examine whether the proposed idea for RM of multiply imaged quasars at the high luminosity and high redshift ends might become realistic in the near future, we estimate the expected number of quasars that would be multiply imaged by galaxy clusters. Given the many uncertainties involved (see for example section 5), our goal is merely to obtain an order-of-magnitude estimate. We focus on the observational properties of three observatories: Euclid, the Roman Space Telescope, and JWST. The basic idea is to calculate for each observatory, i.e., a filter set and depth limit, how many quasars in the relevant BH mass and redshift regimes fall within the caustics of lensing clusters so that they would be multiply imaged, and would be bright enough, folding in the magnification, to be detected. We adopt relevant quasar luminosity functions (QLFs) and generally follow known prescriptions for the lensing optical depth, where we assume an explicit cluster mass function – from which an Einstein radius function is calculated – and then calculate the magnification bias and expected fraction (and total number) of multiply imaged quasars in the relevant cluster mass and redshift ranges.
@@ 3.2 The total expected number of quasars
For the Euclid survey we focus on the detection of typical ”blue” quasars and adopt the spectrum from Vanden Berk et al. (2001). Following the Euclid overview paper (Laureijs et al., 2011), we primarily use the QLF from Willott et al. (2010), but also for comparison employ the Harikane et al. (2023) one. For each redshift bin we choose as the selection band the band in which the rest frame Å falls. If the redshifted wavelength is lower (or higher) than the survey’s available bands, we choose the nearest band to it. Since in this work we are particularly interested in lensed high-mass and high-redshift quasars, we perform the calculation for those two regimes specifically.
We follow the same procedure for the Nancy Grace Roman Space Telescope. Both observatories cover roughly similar wavelength ranges into the near-infrared. While the nominal FOV for Roman is about half of that of Euclid, Roman has a two times in diameter larger mirror and the observational limits are deeper than the ones for Euclid, hence we may expect higher numbers of quasars to be detectable with it. Again our focus remains on the high-redshift and high-luminosity quasars and as above we adopt the typical SED from Vanden Berk et al. (2001).

### v2

@@ Abstract
Mass estimates of black holes (BHs) in the centers of Active Galactic Nuclei (AGN) often rely on the radius-luminosity relation. However, this relation, usually probed by reverberation mapping (RM), is poorly constrained in the high-luminosity and high-redshift ends due to the very long expected RM lag times. Multiply imaged AGN may offer a unique opportunity to explore the radius-luminosity relation at these ends. In addition to comprising several magnified images enabling a more efficient light-curve sampling, the time delay between multiple images of strongly lensed quasars can also aid in making such RM measurements feasible on reasonable timescales: If the strong-lensing time delay is, for example, of the order of the expected RM time lag, changes in the emission lines in the leading image can be observed around the same time as the changes in the continuum in the trailing image. In this work we probe the typical time-delay distribution in galaxy-cluster lenses and estimate the number of both high-mass ( M⊙), and high-redshift () quasars that are expected to be strongly lensed by clusters. We find that up to several dozen thousand M- M⊙ broad-line AGN at should be multiply imaged by galaxy clusters and detectable with JWST, hundreds with Euclid and several thousands with the Roman Space Telescope, across the whole sky. These could supply an important calibration for the BH mass scaling in the early Universe.
@@ 1 Introduction
Super massive black holes (SMBHs) are found in the centers of most galaxies. The masses of those SMBHs are of much interest as their evolution is not well understood, and since they are believed to play a major role in their host galaxy evolution. In fact, tight relations seem to hold between the mass of a SMBH and the properties of its host galaxy such as the luminosity or the velocity dispersion of the bulge (e.g., Ferrarese & Merritt, 2000; Kormendy & Ho, 2013; Reines & Volonteri, 2015; Ding et al., 2020). Measuring SMBH masses as a function of redshift can also shed light onto black hole (BH) formation and the physics of accretion processes (e.g., Wang et al., 2021; Volonteri, 2012), and their mass evolution through cosmological time might help constrain different cosmological models (e.g., Fanidakis et al., 2012; Habouzit et al., 2022).
@@
Unfortunately, RM of high-luminosity and high-redshift quasars faces several difficulties. High-luminosity quasars are presumed to have large BLR sizes, resulting in long RM time lags between continuum and emission-line variations. Similarly, high-redshift quasars and AGN (e.g., Wang et al., 2021; Larson et al., 2023; Goulding et al., 2023; Furtak et al., 2024) will also show longer observed RM time lags compared to their lower-redshift counterparts, due to cosmological time dilation. As a result, observations for several decades are needed for those RM campaigns (Kaspi, 2018, 2021). Nevertheless, several attempts have been recently made to conduct RM for those types of AGN. In one study, Kaspi et al. (2021) observed photometrically 11 quasars at redshifts and luminosities for almost twenty years while six of them were monitored spectrophotometrically for 13 years. Their study manages to significantly measure RM time lags of the C iv Å doublet in three objects and a C iii] Å RM time lag in one object. Scaling the BLR size with the UV continuum luminosity they conclude that their objects lie on the previously measured relation. Shen et al. (2023) preformed RM using 11-year photometric and a 7-year spectroscopic data for 849 broad-line quasars. They report 339 significantly measured RM time lag detections for the different emission lines. Although they claim that on average the R-L relations follow existing ones, they report an intrinsic scatter of dex in the lags for the H line at 4861 Å and the Mg ii] Å doublet, whereas the C iv R-L relation has an even more substantial scatter of dex. It is thus becoming important to understand whether these large scatters stem from intrinsic physical processes at high luminosities (Du et al., 2018a; Czerny et al., 2019; Fonseca Alvarez et al., 2020) or due to the fact that there is often a difficulty to significantly measure RM time lags for such objects given the monitoring times needed.
@@
In this work, we discuss the possibility to probe the R-L relation for the high-luminosity or high-redshift regimes using multiply imaged quasars. If the strong-lensing (SL) time delays (TD) between different multiple images of a lensed quasar are of the same order as the RM time lags, the expected changes in the emission lines can be observed in a leading multiple image around the same time as the changes in the continuum from a trailing multiple image, so that RM of these types of objects can be done with a reasonable and shorter monitoring time-span than for similar, un-lensed quasars. Although the monitoring time will be nominally shorter by one RM time lag, we also discuss here that in practice the gain will be larger, as a lensed quasar will be lensed into several magnified images, supplying a more efficient sampling of the light curve for RM. It is worth mentioning that high quality light curves are also an essential part of velocity-resolved RM (e.g., Bentz et al., 2008; Pancoast et al., 2014; Du et al., 2018b). This is a powerful technique aiming to constrain the kinematics and geometry of the BLR through measuring the RM time lags at various velocities, not only the mean RM time lag as in classical RM. That means that the signal-to-noise and cadence need to be sufficiently high, which might be applicable due to the mentioned lensing effects. In particular, as for the classical RM, at high redshifts the process becomes even more difficult, resulting in the lack of any successful velocity-resolved measurements in this regime.
@@
About a handful of quasars and AGN, mostly around cosmic noon, are known to be multiply imaged by galaxy clusters (e.g., Inada et al., 2003; Oguri et al., 2013; Sharon et al., 2017; Acebron et al., 2022; Furtak et al., 2023a; Napier et al., 2023a), whereas some have also been used for cosmography (e.g., Napier et al., 2023b). Notably, a successful RM of a strongly lensed quasar has been recently done by Williams et al. (2021) where they exploit the TD between the multiple images of the strongly lensed quasar SDSS J2222+2745 to construct a detailed light curve spending over 6 years in the observed frame with 4.5 years of observations with high cadence. They significantly detect a RM time lag of rest-frame days for the C iv emission line, whereas the TD from the two other images they use are of and days. The quasar SDSS J2222+2745 ranged in continuum luminosity with the average value of log10[/erg s-1] = . In this illustrative case the RM time lag was short enough so that the gain from image multiplicity was mainly manifested in additional data points on the light curve, thus importantly allowing to shorten the overall monitoring time. Here, we wish to focus on objects at the high luminosity end, namely continuum UV luminosities erg s-1 (BH masses of log10M), or at high redshifts (i.e., ), for which we expect RM time lags of days, depending on redshift. For example, a redshift , M⊙ quasar (e.g. Wang et al., 2021) would be expected to show an observed RM time lag of about years, requiring decades of monitoring to measure it without lensing multiplicity. In this paper we suggest that SL TDs of lensed quasars may be a useful route to constraining the R-L relation at those regimes. This is becoming especially timely now that JWST has been uncovering a new abundant population of Type I AGN at high redshifts (e.g., Kocevski et al., 2023a; Harikane et al., 2023; Matthee et al., 2023; Larson et al., 2023; Greene et al., 2023), including the first example of a cluster-lensed multiply imaged one at 7 (Furtak et al., 2023b, 2024).
@@
The work is constructed as follows. In section 2 we start by estimating the expected RM time lags for high-luminosity and high-redshift quasars (§2.1), and then estimate the distribution of TDs between the multiple images of lensed systems in typical lensing clusters (§2.2). In section 3 we estimate the expected number of lensed quasars and AGN that could be found with e.g., the JWST, Euclid, or with the Roman Space Telescope, in the near future. The results of this calculation are given in section 4. In section 5 we demonstrate the suggested idea using simulated light curves and discuss its applicability, and in section 6 we conclude the work. Throughout we assume a standard flat CDM cosmology with , , and . Magnitudes are quoted in the AB system (Oke & Gunn, 1983) and all quoted uncertainties represent ranges unless otherwise stated.
@@
In this work we use the terms “quasars” and “AGN” synonymously and whenever we discuss lensed AGN in this work, we explicitly refer to multiply imaged examples. In addition, in this work “time delay” always refers to the gravitational lensing time delay, and for RM time lag we always use the term “time lag”.
2 The distributions of RM time lags and lensing time delays
2.1 RM of quasars: the distributions of RM time lags
@@
As was noted above, RM of quasars relies on the ability to detect temporal variations in the continuum emission from a quasar that lead to, after a time lag which depends on the BLR size, similar variations in the emission lines from the BLR. In this section, we estimate the BLR sizes – and thus RM time lags – of a representative sample of quasars, namely the SDSS quasar catalog, which is based on Data Release 7 (DR7) (Shen et al., 2011) and is available online. This catalog supplies continuum luminosities for either or , depending on redshift.
@@
The distributions of BLR sizes are shown in Fig.1. The expected H RM time-lag distribution is obtained assuming the R-L relation from Bentz et al. (2013), for those quasars that have a measurement of the continuum luminosity at 5100 Å. The expected C iv RM time-lag distribution is obtained assuming the R-L relation from Kaspi et al. (2021) (see their eq. 1), using the continuum UV luminosity at 1350 Å. Note that the C iv subsample consists of 52,427 objects, which have a higher redshift (median redshift of 1.96) and are typically more luminous (median Å erg s-1) than the H subsample (23,093 objects, median redshift of 0.59, median Å erg s-1), the redshift distribution of both subsamples is shown in Fig.2. As can be seen in Fig. 1, the BLR size distributions, plotted as log10(R) where R is in light days, are approximately Gaussian, with a mean and width of = and = for H and C iv, respectively. About %, roughly, of the quasars show C iv RM time lags of days or longer in the rest-frame, and about % of the quasars show H RM time lags of days or longer in the rest-frame. This translates to days for redshifts of , respectively, which we define here as the high redshift regime. Since we explore the typical RM time lags in the rest frame, we should note the intrinsic redshift distribution of the sample. For all AGN are at redshift (i.e. in a relatively narrow redshift range, see Fig. 2). For CIV, are between redshifts and , and the rest are in the tail (extending to redshift ). When we construct the CIV BLR size distribution from only those that are at , we get a similar shape distribution with the mean value shifted slightly to lower values (by ). AGN at redshift, for example, constitute of the total population used to create the CIV BLR size distribution and when using only them, the shape of the distribution still holds, although the mean value of the BLR size distribution is shifted to higher values by . This shift is most likely due to the fact that at higher redshifts, for a given observational depth, only more luminous AGN could be observed. These deviations are thus small enough, allowing us to assume in the following subsection 2.2 and upcoming Fig.3 that the distributions in Fig.1 are representative for a population of AGN at any redshift bin. Note that this procedure does not take into account a possible evolution of the BLR size with redshift (which may not be very significant e.g., Du et al. 2018b), nor other observational biases.
@@ 2.2 The distribution of strong-lensing time delays
In this subsection we aim to obtain a general distribution of the TDs caused by strong lensing in clusters. Galaxy clusters act as gravitational lenses, causing the bending of light. When the projected mass density is larger than some critical density for lensing, strong lensing will occur comprising multiple images of background sources. Using the thin lens approximation, light that travels from a source to the observer will be delayed by some time (Schneider, 1985):
@@
In order to create the TD distribution, we randomly plant sources in the source plane and re-lens them to the image plane using the deflection fields from the model. Typically we use a 100 multiply imaged sources in each case. We calculate the TDs between all available pairs using eq. 1 and repeat this process for different source redshifts, in the ranges of , , , , and . The SL time-delay distributions are plotted in Fig.3.
@@
In Fig. 3 we also compare the lensing TD probability distributions to the RM time lags expected from the BLR sizes, recalculated to match each redshift bin ,as we assume that this distribution is representative at all redshifts, which is a sufficient approximation for our purpose, namely showing that the RM time lag distribution is of the order of TDs by galaxy clusters. More on this assumption in previous subsection 2.1. We see that there is an overlap in the distributions such that TDs by both types of galaxy clusters cover the entire range of expected RM time lags, with a higher probability of having the relevant TDs, i.e., of order the RM time lag, for higher-luminosity and higher-redshift quasars. About of the TD pairs in all clusters overlap with the range of RM time lags, depending on the redshift of the lensed source. We see that the peak of the TD distribution caused by lower mass clusters, which should be more abundant, agrees better with the peak of the RM time lag distributions, but those clusters would have also less multiple images per cluster in proportion to their smaller critical area size (which is roughly proportional to the square root of the mass), such that both populations should be useful. Also overplotted is a line marking a typical TD expected by a massive galaxy lens – where we nominally adopt for this calculation a SIS galaxy of km s-1 placed at redshift 0.5 (Oguri, 2019), demonstrating that TD in the case of lens galaxies are too short for a significant gain in the wait time needed for RM of high-redshift, high-mass quasars, and that galaxy clusters are needed for that purpose.
@@ 3 Expected numbers of lensed quasars
To examine whether the proposed idea for RM of multiply imaged quasars at the high luminosity and high redshift ends might become realistic in the near future, we estimate the expected number of quasars that would be multiply imaged by galaxy clusters. Given the many uncertainties involved (see for example subsection 5.2), our goal is merely to obtain an order-of-magnitude estimate. We focus on the observational properties of three observatories: Euclid, the Roman Space Telescope, and JWST. Euclid is an optical and NIR space telescope with a mirror of size 1.2 meters in diameter, covering the wavelengths between 550 nm and 2 m. Covering the same wavelength range, the soon to be launched Roman Space Telescope will have a 2.4 m primary mirror and will achieve greater depth with the same observation time. Lastly, JWST is mainly a NIR space telescope with a 6.5m mirror, with wavelength coverage ranging from 0.6 m to 28.3 m. The basic idea is to calculate for each observatory, i.e., a filter set and depth limit, how many quasars in the relevant BH mass and redshift regimes fall within the caustics of lensing clusters so that they would be multiply imaged, and would be bright enough, folding in the magnification, to be detected. We adopt relevant quasar luminosity functions (QLFs) and generally follow known prescriptions for the lensing optical depth, where we assume an explicit cluster mass function – from which an Einstein radius function is calculated – and then calculate the magnification bias and expected fraction (and total number) of multiply imaged quasars in the relevant cluster mass and redshift ranges.
@@ 3.2 The total expected number of quasars
For the Euclid survey we focus on the detection of typical ”blue” quasars and adopt the spectrum from Vanden Berk et al. (2001). Following the Euclid overview paper (Laureijs et al., 2011), we primarily use the QLF from Willott et al. (2010), but also for comparison employ the Harikane et al. (2023) one. The observational depth for the different bands is also adopted from Laureijs et al. 2011. For each redshift bin we choose as the selection band the band in which the rest frame Å falls. If the redshifted wavelength is lower (or higher) than the survey’s available bands, we choose the nearest band to it. Since in this work we are particularly interested in lensed high-mass and high-redshift quasars, we perform the calculation for those two regimes specifically.
We follow the same procedure for the Nancy Grace Roman Space Telescope. Both observatories cover roughly similar wavelength ranges into the near-infrared. While the nominal FOV for Roman is about half of that of Euclid, Roman has a two times in diameter larger mirror and the observational limits are deeper than the ones for Euclid, hence we may expect higher numbers of quasars to be detectable with it. We adopt the depth values for the different bands from the Roman Space Telescope website 111https://www.stsci.edu/files/live/sites/www/files/home/roman/_documents/roman-science-sheet.pdf. Again our focus remains on the high-redshift and high-luminosity quasars and as above we adopt the typical SED from Vanden Berk et al. (2001).